# Long-Horizon Model-Based Offline Reinforcement Learning Without Conservatism

### v1

@@ Abstract
Popular offline reinforcement learning (RL) methods rely on conservatism, either by penalizing out-of-dataset actions or by restricting rollout horizons. In this work, we question the universality of this principle and instead revisit a complementary one: a Bayesian perspective. Rather than enforcing conservatism, the Bayesian approach tackles epistemic uncertainty in offline data by modeling a posterior distribution over plausible world models and training a history-dependent agent to maximize expected rewards, enabling test-time generalization. We first illustrate, in a bandit setting, that Bayesianism excels on low-quality datasets where conservatism fails. We then scale this principle to realistic tasks, identifying key design choices, such as layer normalization in the world model and adaptive long-horizon planning, that mitigate compounding error and value overestimation. These yield our practical algorithm, Neubay, grounded in the neutral Bayesian principle. On D4RL and NeoRL benchmarks, Neubay generally matches or surpasses leading conservative algorithms, achieving new state-of-the-art on 7 datasets. Notably, it succeeds with rollout horizons of several hundred steps, contrary to dominant practice. Finally, we characterize datasets by quality and coverage, showing when Neubay is preferable to conservative methods. Together, we argue Neubay lays the foundation for a new practical direction in offline and model-based RL. [^1] Code is available at https://github.com/twni2016/neubay.
@@ 1 Introduction
However, scaling the Bayesian principle to realistic tasks is challenging, as it requires solving an approximate epistemic POMDP. We identify three key challenges: (1) compounding error in world models, where inaccuracies grow rapidly with horizon (Lambert et al., 2022); (2) value overestimation, since Bayesian RL lacks explicit penalties for out-of-dataset actions and is vulnerable to extrapolation error (Fujimoto et al., 2019); and (3) training agents with long-term memory to enable test-time adaptation (Ni et al., 2023). Each aligns with a major research area: model-based RL, offline RL, and partially observable RL. This helps explain why prior Bayesian-inspired algorithms often reintroduce conservatism, through uncertainty penalties, short horizons, or reductions to model-free RL (Ghosh et al., 2022; Chen et al., 2021c; Jeong et al., 2023).
We build a practical algorithm, Neubay, to address these challenges. First, we show that layer normalization (Ba et al., 2016), effective in model-free RL (Ball et al., 2023), also helps world models mitigate compounding error. Second, we observe that model-generated rewards are less biased than bootstrapped values, allowing Bayesian RL to amortize overestimation risk through long-horizon planning. To keep such rollouts reliable, we truncate them using epistemic uncertainty as a threshold (Frauenknecht et al., 2024; Zhan et al., 2021), an alternative use of uncertainty beyond penalties. Finally, we leverage advances in recurrent RL (Morad et al., 2024; Luo et al., 2024a) to stabilize training and support long-term memory. Together, these components adapt insights from online RL to make Bayesian offline RL practical.
@@ 2 Background on Offline RL
$\max_{\pi}J(\pi,m^{*})\mathrel{\mathop{:}}=\mathbb{E}{\left[\sum\nolimits_{t=0}^{\infty}\gamma^{t}r_{t+1}\,\middle|\,s_{0}\sim\rho,a_{t}\sim\pi(h_{t}),(r_{t+1},s_{t+1})\sim m^{*}(s_{t},a_{t})\right]}.$ (1)
@@
$\max_{\pi}J(\pi;\mathfrak{M}_{\mathcal{D}}):=\max_{\pi}\min_{m\in\mathfrak{M}_{\mathcal{D}}}J(\pi,m).$ (3)
@@
$\max_{\pi}J(\pi;\mathbb{P}_{\mathcal{D}})\mathrel{\mathop{:}}=\max_{\pi}\mathbb{E}_{m\sim\mathbb{P}_{\mathcal{D}}}{\left[J(\pi,m)\right]}.$ (4)
@@ 3 Illustrative Example: When Is Bayesianism Better?
This skewed dataset $\mathcal{D}$ leaves the true reward parameter $p^{*}_{1}$ for arm $1$ completely unobserved, inducing substantial epistemic uncertainty on $p^{*}_{1}$. Theoretically, under an uninformative prior for Bernoulli rewards, this uncertainty corresponds to the entropy of $\mathrm{Unif}[0,1]$. In practice, we approximate the posterior by fitting a reward-model ensemble using Gaussian outputs (Chua et al., 2018). As shown in Figure 2, after pretraining on $\mathcal{D}$, the ensemble predictions concentrate around the observed arm 0 but sharply disagree on the unobserved arm $1$, with estimated uncertainty about $10\times$ larger than on arm 0. Next, we test the Bayesian method from Equation 4, where planning begins at $t=0$ and truncates at $t=T$. Dynamics are not modeled due to the bandit structure, and there is no concern about compounding error.
@@
Uncertainty penalty hurts generalization, but Bayesian agents adapt to worse and better cases. We study conservatism by adding an uncertainty penalty to the Bayesian objective. Such a design is common in empirical algorithms (Yu et al., 2020), and can be viewed as risk-sensitive Bayesian RL (Rigter et al., 2021) or a soft robust MDP (Zhang et al., 2024b) (see Appendix B). Since the estimated uncertainty on arm $1$ is much larger than arm 0, a sufficiently large penalty coefficient $\lambda$ forces the resulting algorithm to always select arm 0, regardless of the test-time environment.
@@ 4 Neubay: A Practical Bayesian-Principled Algorithm
The Bayesian objective (Equation 4) is conceptually simple and, as seen in the bandit example, clearly outperforms conservatism on low-quality data. Extending it to MDPs raises three challenges: compounding model error, value overestimation, and instability in training history-dependent agents. We address these with ensembles for posterior approximation (subsection 4.1), adaptive long-horizon planning (subsection 4.2), and stabilized recurrent training (subsection 4.3). Together these yield our practical algorithm, Neubay (subsection 4.4), whose design choices we analyze in subsection 5.2.
4.1 Controlling Compounding Errors in Epistemic POMDP Modeling
A prerequisite for optimizing Equation 4 is to approximate the epistemic POMDP $\mathbb{P}_{\mathcal{D}}$ with a learned posterior $\hat{\mathbb{P}}_{\mathcal{D}}$. Following popular model-based RL methods (Chua et al., 2018; Yu et al., 2020), we model the posterior with deep ensembles (Lakshminarayanan et al., 2017), a simple method that balances predictive accuracy with uncertainty quantification (Ovadia et al., 2019; Gustafsson et al., 2020). Our deep ensemble is a set of neural networks $\mathbf{m}_{\boldsymbol{\theta}}=\{m_{\theta^{n}}\}_{n=1}^{N}$ where each model outputs a Gaussian distribution over next state $s^{\prime}$ and reward $r$: $m_{\theta^{n}}(s,a)=\mathcal{N}(\mu_{\theta^{n}}(s,a),\sigma_{\theta^{n}}(s,a))$, with parameters $\boldsymbol{\theta}=\{\theta^{n}\}_{n=1}^{N}$ independently initialized at random. Each $m_{\theta^{n}}$ is trained on the same dataset using the maximum likelihood (MLE) loss and evaluated by mean squared error (MSE) on a shared validation set (Yu et al., 2020). Deep ensembles provide a practical approximation to Bayesian neural networks, as they capture diverse modes of the Bayesian posterior (Fort et al., 2019; Wilson & Izmailov, 2020). The induced posterior is $\hat{\mathbb{P}}_{\mathcal{D}}(m)=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}(m=m_{\theta^{n}})$ forms a uniform distribution over ensemble members. Epistemic uncertainty can be estimated as the standard deviation of the mean predictions (Lakshminarayanan et al., 2017) of the ensemble members: $U_{\boldsymbol{\theta}}(s,a)=\mathrm{std}(\{\mu_{\theta^{n}}(s,a)\}_{n=1}^{N}).$ This simple design enables direct comparison with popular methods, although multi-step prediction (Lin et al., 2025) and richer uncertainty quantification (Qiao et al., 2025) may further improve performance.
@@ 4.1 Controlling Compounding Errors in Epistemic POMDP Modeling
Design choice 1: larger ensemble size $N$ for posterior fidelity. While deep ensembles approximate Bayesian neural networks, their epistemic fidelity relies on member diversity. With horizons extending hundreds of steps, compounding errors make faithful posteriors essential, rendering small ensembles (e.g., $N=5$ in MBPO (Janner et al., 2019)) inadequate.
@@
Design choice 2: layer normalization in the world model to control compounding errors. A fundamental problem with world models is that small prediction errors compound during multi-step planning (Talvitie, 2014). To mitigate this, we apply layer normalization (LN) (Ba et al., 2016) to each world model $m_{\theta^{n}}$. Similar to its role in reducing value overestimation (Ball et al., 2023), LN mitigates compounding model error. To address this, we structure the world model as a delta predictor and apply layer normalization (LN) (Ba et al., 2016) to its hidden features. Concretely, the model predicts the next state as $\mathbb{E}{\left[\hat{s}^{\prime}\right]}=s+\mathbf{W}^{\top}\mathrm{ReLU}(\mathrm{LN}(\psi(s,a))$, with features $\psi(s,a)\in\mathbb{R}^{k}$ and output weights $\mathbf{W}\in\mathbb{R}^{k\times|\mathcal{S}|}$. For LN without affine parameters, under $\ell_{2}$ norm:
@@
4.2 Why and How Do we Perform Long-Horizon Planning?
Where to start planning? From the Bayesian objective (Equation 4), it is natural to initiate planning rollouts by sampling initial states $s_{0}\sim\rho_{\mathcal{D}}$, where $\rho_{\mathcal{D}}$ is the empirical initial-state distribution. But this underrepresents states appearing later in real trajectories, as compounding errors make them harder to reach (Lambert et al., 2022; Lin et al., 2025). While layer normalization substantially mitigates this issue, the errors at later time steps can still grow exponentially (e.g., Figure 5). To maintain stability, we sample starting states $s_{t}\sim\mathcal{D}$ from any timestep $t$, following MBPO-style branched rollouts (Janner et al., 2019). Since the environment appears as an epistemic POMDP to the agent, we also provide the corresponding history $h_{t}=(s_{0:t},a_{0:t-1},r_{1:t})\in\mathcal{D}$, which we refer to as the initial history.
Why do we need long-horizon planning? Ideally, under a correct posterior, the Bayesian principle calls for full-horizon planning. In practice, however, compounding errors make full-horizon rollouts unreliable. Even so, long-horizon rollouts remain valuable when they are informative for augmenting training data (Young et al., 2023). Beyond augmentation and exploration, we identify a new role for Bayesian offline RL: long-horizon rollouts actively help mitigate value overestimation.
To illustrate, starting from a real history $h_{t}\in\mathcal{D}$, we sample $m_{\theta}\sim\mathbf{m}_{\boldsymbol{\theta}}$ and generate a trajectory of length $H$ where $\hat{a}_{t+j}=\pi(\hat{h}_{t+j})$ comes from a deterministic policy, $(\hat{r}_{t+j+1},\hat{s}_{t+j+1})\sim m_{\theta}(\hat{s}_{t+j},\hat{a}_{t+j})$, and $\hat{h}_{t}=h_{t}$. Applying one-step Bellman backups on the Bayesian value function along this rollout, i.e., $Q^{\text{Bayes}}(\hat{h}_{t+j},\hat{a}_{t+j})\leftarrow\hat{r}_{t+j+1}+\gamma Q^{\text{Bayes}}(\hat{h}_{t+j+1},\pi(\hat{h}_{t+j+1}))$, $0\leq j<H$, we obtain an approximate $H$-step backup on the real history:
This decomposition highlights the bias trade-off: imagined rewards can be low-bias if the model generalizes, while the bootstrapped term – more susceptible to overestimation (Kumar et al., 2019; Sims et al., 2024) – is exponentially discounted with $H$. We formalize this effect in subsection C.2, extending the analysis of Sims et al. (2024). While prior works in offline RL mitigate overestimation through conservatism (e.g., uncertainty penalties), which hurts generalization, Neubay instead absorbs the overestimation risk across long horizons without conservatism.
How should we truncate rollouts? Given that Bayesian RL favors long-horizon planning, a key question is not simply when to truncate, but where, since model errors depend on specific $(s,a)$ pairs. A natural criterion is the model’s uncertainty estimate $U_{\boldsymbol{\theta}}(s,a)$, which correlates with prediction error. Prior work has pursued this conservatively, combining uncertainty threshold with short horizon caps (Pan et al., 2020; Zhan et al., 2021; Zhang et al., 2023; Frauenknecht et al., 2024). In contrast, we remove restriction on the maximum horizon.
Epistemic uncertainty is highly non-uniform even within $\mathcal{D}$: frequently visited $(s,a)$ pairs yield low uncertainty, while rarely seen ones result in high uncertainty. This reflects that epistemic uncertainty concerns the true model $m^{*}$, not the empirical model $m_{\mathcal{D}}$, which can be formalized via standard concentration bounds in the frequentist sense. [^3] With probability $1-\delta$, for $(s,a)\in\mathrm{supp}_{\mathcal{S}\times\mathcal{A}}(\mathcal{D})$, $\textsc{tv}(m_{\mathcal{D}}(s,a),m^{*}(s,a))\leq c_{m^{*},\delta}/\sqrt{n_{\mathcal{D}}(s,a)}$, where $n_{\mathcal{D}}(s,a)$ is the visitation count in $\mathcal{D}$ and $c_{m^{*},\delta}$ is a constant (Kumar et al., 2020, Section D.3). As shown in Figure 4, most empirical CDFs are long-tailed, with dataset-dependent skewness reflecting the underlying visitation distribution. This motivates our use of data-dependent truncation thresholds, described next.
Design choice 3: uncertainty threshold $\mathcal{U}(\zeta)$ for rollout truncation. Let $\zeta\in[0,1]$, we define
4.3 Stable Recurrent Off-Policy RL Training
The last core component of our algorithm is RL training. We follow the MBPO paradigm (Janner et al., 2019), using model-free RL on a mixture of real and model-generated data. Since the real data is off-policy and $\hat{\mathbb{P}}_{\mathcal{D}}$ induces a POMDP for the agent, we adopt recurrent off-policy RL (Ni et al., 2022) rather than on-policy RL (Fellows et al., 2025). We use a recurrent actor $\pi_{\nu}(a_{t}\mid h_{t})$ and a recurrent critic $Q_{\omega}(h_{t},a_{t})$, each with a separate RNN encoder ($\nu_{\phi}(h_{t})$ and $\omega_{\phi}(h_{t})$, respectively) for stability (Ni et al., 2022). Long rollout horizons and the need for test-time generalization require memory spanning entire episodes (up to $1000$ steps in our tasks), which exceeds the capacity of vanilla RNNs (Ni et al., 2023). To address this, we extend memoroid (Morad et al., 2024) to actor-critic architectures using linear recurrent units (LRUs) (Orvieto et al., 2023) as the RNN encoders, enabling both long-term memory and parallel optimization. While a memoroid-based architecture is powerful, it was originally designed for online POMDPs. Bridging the gap to epistemic POMDPs in offline RL introduces new challenges, motivating several design choices that we detail below.
Design choice 5: small context encoder learning rate $\eta_{\phi}$ for training stability. The learning rate is a critical hyperparameter in deep RL, and recurrent off-policy RL is especially sensitive to it. In online POMDPs, RESeL (Luo et al., 2024a) shows that recurrent encoders require much smaller (around $30\times$) learning rates than those of actor or critic heads, since history representations can diverge exponentially with history length even under tiny parameter changes.
@@ 4.3 Stable Recurrent Off-Policy RL Training
5:   Sample $m_{\theta}\sim\mathbf{m}_{\boldsymbol{\theta}}$

### v2

@@ Abstract
Popular offline reinforcement learning (RL) methods rely on conservatism, either by penalizing out-of-dataset actions or by restricting rollout horizons. In this work, we question the universality of this principle and instead revisit a complementary one: a Bayesian perspective. Rather than enforcing conservatism, the Bayesian approach tackles epistemic uncertainty in offline data by modeling a posterior distribution over plausible world models and training a history-dependent agent to maximize expected rewards, enabling test-time generalization. We first illustrate, in a bandit setting, that Bayesianism excels on low-quality datasets where conservatism fails. We then scale this principle to realistic tasks and show that long-horizon planning is critical for reducing value overestimation once conservatism is removed. To make this feasible, we introduce key design choices for performing and learning from long-horizon rollouts while controlling compounding errors. These yield our algorithm, Neubay, grounded in the neutral Bayesian principle. On D4RL and NeoRL benchmarks, Neubay generally matches or surpasses leading conservative algorithms, achieving new state-of-the-art on 7 datasets. Notably, it succeeds with rollout horizons of several hundred steps, contrary to dominant practice. Finally, we characterize datasets by quality and coverage, showing when Neubay is preferable to conservative methods. Together, we argue Neubay lays the foundation for a new practical direction in offline and model-based RL. [^1] Code is available at https://github.com/twni2016/neubay.
@@ 1 Introduction
However, scaling the Bayesian principle to realistic tasks is challenging, as it requires solving an approximate epistemic POMDP. We identify three key challenges: (1) value overestimation, which becomes central once Bayesian RL abandons conservatism and removes penalties for out-of-dataset actions (Fujimoto et al., 2019); (2) compounding error in world models, where inaccuracies grow rapidly with horizon (Lambert et al., 2022); and (3) training agents with long-term memory to enable test-time adaptation (Ni et al., 2023). Each aligns with a major research area: offline RL, model-based RL, and partially observable RL. This helps explain why prior Bayesian-inspired algorithms often reintroduce conservatism, through uncertainty penalties, short horizons, or reductions to model-free RL (Ghosh et al., 2022; Chen et al., 2021c; Jeong et al., 2023).
We build a practical algorithm, Neubay, to address these challenges. We first show that long-horizon rollouts can themselves reduce value overestimation once conservatism is removed. We then introduce design choices that make such long-horizon rollouts feasible. To control compounding errors, we truncate rollouts adaptively using epistemic uncertainty as a threshold (Frauenknecht et al., 2024; Zhan et al., 2021), an alternative use of uncertainty beyond penalties, and apply layer normalization (Ba et al., 2016) within the world model. To stabilize learning from long horizons, we leverage recent advances in online recurrent RL with long-term memory (Morad et al., 2024; Luo et al., 2024a).
@@ 2 Background on Offline RL
$\max\nolimits_{\pi}J(\pi,m^{*})\mathrel{\mathop{:}}=\mathbb{E}{\left[\sum\nolimits_{t=0}^{\infty}\gamma^{t}r_{t+1}\,\middle|\,s_{0}\sim\rho,a_{t}\sim\pi(h_{t}),(r_{t+1},s_{t+1})\sim m^{*}(s_{t},a_{t})\right]}.$ (1)
@@
$\max\nolimits_{\pi}J(\pi;\mathfrak{M}_{\mathcal{D}}):=\max\nolimits_{\pi}\min\nolimits_{m\in\mathfrak{M}_{\mathcal{D}}}J(\pi,m).$ (3)
@@
$\max\nolimits_{\pi}J(\pi;\mathbb{P}_{\mathcal{D}})\mathrel{\mathop{:}}=\max\nolimits_{\pi}\mathbb{E}_{m\sim\mathbb{P}_{\mathcal{D}}}{\left[J(\pi,m)\right]}.$ (4)
@@
Approximate posterior and uncertainty quantifier. To instantiate the Bayesian objective (Equation 4), we require a tractable approximation $\hat{\mathbb{P}}_{\mathcal{D}}$ to the posterior over models. Following common practice in model-based RL (Chua et al., 2018; Yu et al., 2020), we approximate $\mathbb{P}_{\mathcal{D}}$ with a finite ensemble of learned models $\mathbf{m}_{\boldsymbol{\theta}}=\{m_{\theta^{n}}\}_{n=1}^{N}$, which induces the empirical posterior $\hat{\mathbb{P}}_{\mathcal{D}}(m)=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}(m=m_{\theta^{n}}).$ The ensemble provides a practical approximation to Bayesian posteriors by capturing diverse modes (Fort et al., 2019; Wilson & Izmailov, 2020). Given a state-action pair $(s,a)$, we adopt the disagreement among the mean predictions (Lakshminarayanan et al., 2017) as uncertainty quantifier, $U_{\boldsymbol{\theta}}(s,a)=\mathrm{std}(\{\mu_{\theta^{n}}(s,a)\}_{n=1}^{N}),$ which is shown to be empirically robust under dataset shift (Ovadia et al., 2019). We defer architectural and training details of the ensemble to Appendix G.
@@ 3 Illustrative Example: When Is Bayesianism Better?
This skewed dataset $\mathcal{D}$ leaves the true reward parameter $p^{*}_{1}$ for arm $1$ completely unobserved, inducing substantial epistemic uncertainty on $p^{*}_{1}$. Theoretically, under an uninformative prior for Bernoulli rewards, this uncertainty corresponds to the entropy of $\mathrm{Unif}[0,1]$. We approximate the posterior by fitting an ensemble of reward models $\mathbf{m}_{\boldsymbol{\theta}}$ using Gaussian outputs. As shown in Figure 2, after pretraining on $\mathcal{D}$, the ensemble predictions concentrate around the observed arm 0 but sharply disagree on the unobserved arm $1$, with estimated uncertainty about $10\times$ larger than on arm 0. Next, we test the Bayesian method from Equation 4, where planning begins at $t=0$ and truncates at $t=T$. Dynamics are not modeled due to the bandit structure, and there is no concern about compounding error.
@@
Uncertainty penalty hurts generalization, but Bayesian agents adapt to worse and better cases. We study conservatism by penalizing the reward with an uncertainty-based term $\lambda\,U_{\boldsymbol{\theta}}(a)$. Such a design is common in empirical algorithms (Yu et al., 2020), and can be viewed as risk-sensitive Bayesian RL (Rigter et al., 2021) or a soft robust MDP (Zhang et al., 2024b) (see Appendix B). Since the estimated uncertainty on arm $1$ is much larger than arm 0 (i.e, $U_{\boldsymbol{\theta}}(1)\gg U_{\boldsymbol{\theta}}(0)$), a sufficiently large coefficient $\lambda$ forces the resulting algorithm to always select arm 0, regardless of the test-time environment.
@@ 4 Neubay: A Practical Bayesian-Principled Algorithm
The Bayesian agent is conceptually simple and, as seen in the bandit example, clearly outperforms conservative agents on low-quality data. Extending this principle from bandits to realistic MDPs, however, is challenging: the full-horizon rollouts implied by the Bayesian objective (Equation 4) can suffer from severe compounding model errors (Luo et al., 2024b; Lin et al., 2025). While this may suggest restricting to short-horizon rollouts, we first show that long-horizon rollouts are in fact necessary for Bayesian agents, as longer horizons reduce value overestimation, a central challenge once conservatism is removed (subsection 4.1). We then introduce a set of design choices for both performing and learning from long-horizon rollouts (subsection 4.2subsection 4.4), yielding our practical algorithm, Neubay (subsection 4.5).
4.1 Why Do We Need Long-Horizon Rollouts?
Ideally, under a correct posterior, the Bayesian principle calls for full-horizon rollout. In practice, however, compounding errors make full-horizon rollouts unreliable. To mitigate this issue, prior work in offline RL typically adopts a combination of short-horizon rollouts and conservatism: short horizons limit compounding errors, while conservatism (e.g., uncertainty penalties) reduces value overestimation due to the extrapolation error (Fujimoto et al., 2019; Kumar et al., 2019).
In contrast, our focus is on Bayesian agents that aim to generalize beyond the dataset, as illustrated by section 3. This requires abandoning conservatism, but in doing so removes an explicit mechanism for controlling value overestimation. We show that, perhaps surprisingly, long-horizon rollouts can themselves serve this role by actively reducing overestimation compared to short-horizon rollouts.
To illustrate, starting from a real history $h_{t}\in\mathcal{D}$, we sample a model $m_{\theta}$ from the ensemble $\mathbf{m}_{\boldsymbol{\theta}}$ and generate a trajectory of length $H$, where $\hat{a}_{t+j}=\pi(\hat{h}_{t+j})$ comes from a deterministic policy, $(\hat{r}_{t+j+1},\hat{s}_{t+j+1})\sim m_{\theta}(\hat{s}_{t+j},\hat{a}_{t+j})$, and $\hat{h}_{t}=h_{t}$. Applying one-step Bellman backups on the Bayesian value function along this rollout, i.e., $Q^{\text{Bayes}}(\hat{h}_{t+j},\hat{a}_{t+j})\leftarrow\hat{r}_{t+j+1}+\gamma Q^{\text{Bayes}}(\hat{h}_{t+j+1},\pi(\hat{h}_{t+j+1}))$, $0\leq j<H$, we obtain an approximate $H$-step backup on the real history:
This decomposition highlights the bias trade-off: imagined rewards can be low-bias if the model generalizes, while the bootstrapped term – more susceptible to overestimation (Kumar et al., 2019; Sims et al., 2024) – is exponentially discounted with $H$. We formalize this effect in subsection C.2, extending the analysis of Sims et al. (2024). Importantly, this role of long-horizon rollouts in mitigating overestimation is distinct from their use for data augmentation or exploration (Young et al., 2023). Rather than relying on conservatism, long horizons absorb overestimation risk by shifting estimation from bootstrapping toward model-generated rewards. The cost of longer horizons is increased compounding error, which we address in subsection 4.2 and subsection 4.3.
4.2 How Do We Perform Long-Horizon Rollouts?
Where to start planning? From the Bayesian objective (Equation 4), it is natural to initiate planning rollouts by sampling initial states $s_{0}\sim\rho_{\mathcal{D}}$, where $\rho_{\mathcal{D}}$ is the empirical initial-state distribution. But this underrepresents states appearing later in real trajectories, as compounding errors make them harder to reach. Following MBPO-style branching mechanism (Janner et al., 2019), we sample starting states $s_{t}\sim\mathcal{D}$ from any timestep $t$ to maintain coverage. Since the environment appears as an epistemic POMDP to the agent, we also provide the corresponding history $h_{t}=(s_{0:t},a_{0:t-1},r_{1:t})\in\mathcal{D}$, which we refer to as the initial history.
How to truncate long-horizon rollouts? Since model errors depend on specific $(s,a)$ pairs, a key question is not simply when to truncate rollouts, but where. A natural criterion is the model’s uncertainty estimate $U_{\boldsymbol{\theta}}(s,a)$, which correlates with prediction error. In practice, epistemic uncertainty is highly non-uniform even within $\mathcal{D}$: frequently visited $(s,a)$ pairs yield low uncertainty, while rarely seen ones result in high uncertainty. This reflects that epistemic uncertainty concerns the true model $m^{*}$, not the empirical model $m_{\mathcal{D}}$, which can be formalized via standard concentration bounds in the frequentist sense. [^3] With probability $1-\delta$, for $(s,a)\in\mathrm{supp}_{\mathcal{S}\times\mathcal{A}}(\mathcal{D})$, $\textsc{tv}(m_{\mathcal{D}}(s,a),m^{*}(s,a))\leq c_{m^{*},\delta}/\sqrt{n_{\mathcal{D}}(s,a)}$, where $n_{\mathcal{D}}(s,a)$ is the visitation count in $\mathcal{D}$ and $c_{m^{*},\delta}$ is a constant (Kumar et al., 2020, Section D.3). As shown in Figure 4, most empirical CDFs are long-tailed, with dataset-dependent skewness reflecting the underlying visitation distribution. Prior work has adopted uncertainty threshold conservatively, combining it with short horizon caps (Pan et al., 2020; Zhan et al., 2021; Zhang et al., 2023; Frauenknecht et al., 2024). In contrast, we remove any restriction on the maximum horizon and let uncertainty alone determine truncation.
Design choice 1: uncertainty threshold $\mathcal{U}(\zeta)$ for rollout truncation. Let $\zeta\in[0,1]$, we define
4.3 World Model Architecture for Long-Horizon Rollouts
To support long-horizon rollouts, we introduce two simple and effective architectural choices for the world model ensemble that mitigate compounding errors and improve posterior fidelity.
@@ 4.3 World Model Architecture for Long-Horizon Rollouts
Design choice 2: larger ensemble size $N$. While deep ensembles approximate Bayesian posteriors, their fidelity relies on member diversity. Under long-horizon rollouts, compounding errors amplify posterior inaccuracies, rendering small ensembles (e.g., $N=5$ in MBPO (Janner et al., 2019)) inadequate. In our main experiments, we therefore use a larger ensemble size ($N=100$), which remains computationally feasible due to parallelization.
@@
Design choice 3: layer normalization in the world model. We further find that applying layer normalization (LN) (Ba et al., 2016) within the world model significantly mitigates compounding errors for long-horizon rollouts. This mirrors its role in reducing extrapolation error in model-free RL (Ball et al., 2023). Concretely, we structure the world model as a delta predictor and apply LN to its hidden features. The model predicts the next state as $\mathbb{E}{\left[\hat{s}^{\prime}\right]}=s+\mathbf{W}^{\top}\mathrm{ReLU}(\mathrm{LN}(\psi(s,a))$, with features $\psi(s,a)\in\mathbb{R}^{k}$ and output weights $\mathbf{W}\in\mathbb{R}^{k\times|\mathcal{S}|}$. For LN without affine parameters, under $\ell_{2}$ norm:
@@
5:   Sample $m_{\theta}\sim\mathbf{m}_{\boldsymbol{\theta}}$, fixed for the rollout
@@
4.4 Stable Recurrent Off-Policy RL Using Long-Horizon Rollouts
Lastly, we describe how we train policies from long-horizon rollout data. We follow the MBPO paradigm (Janner et al., 2019), using model-free RL on a mixture of real and model-generated data. Since the real data is off-policy and $\hat{\mathbb{P}}_{\mathcal{D}}$ induces a POMDP for the agent, we adopt recurrent off-policy RL (Ni et al., 2022) rather than on-policy RL (Fellows et al., 2025). We use a recurrent actor $\pi_{\nu}(a_{t}\mid h_{t})$ and a recurrent critic $Q_{\omega}(h_{t},a_{t})$, each with a separate RNN encoder ($\nu_{\phi}(h_{t})$ and $\omega_{\phi}(h_{t})$, respectively) for stability (Ni et al., 2022). Long rollout horizons and the need for test-time generalization require memory spanning entire episodes (up to $1000$ steps in our tasks), which exceeds the capacity of vanilla RNNs (Ni et al., 2023). To address this, we extend memoroid (Morad et al., 2024) to actor-critic architectures using linear recurrent units (LRUs) (Orvieto et al., 2023) as the RNN encoders, enabling both long-term memory and parallel optimization. While a memoroid-based architecture is powerful, it was originally designed for online POMDPs. Bridging the gap to epistemic POMDPs in offline RL introduces new challenges, motivating several design choices that we detail below.
Design choice 5: small context encoder learning rate $\eta_{\phi}$. The learning rate is a critical hyperparameter for training stability, especially in recurrent off-policy RL with long memory. In online POMDPs, RESeL (Luo et al., 2024a) shows that recurrent encoders require much smaller (around $30\times$) learning rates than those of actor or critic heads, since history representations can diverge exponentially with history length even under tiny parameter changes.